OpenStudy (anonymous):

write the equation of the line that has slope -2 and passes through (-4,-5)

OpenStudy (alexandervonhumboldt2):

ok so first y=ax+b where a is slope and b is y-intercept. 1st substitue -2 for a. y=-2a+b. now substitue (x, y) corrdinates and solve for b: -5=-4*(-2)+b. b=-5-8=-13. thus the equationof line is \[y=-2x -13\]